OpenStudy (anonymous):

An architect is planning a rectangular courtyard. The length must be exactly 31 meters, and the perimeter can be at most 100 meters. What values are allowed for the rectangular courtyard's width? Use w for the width (in meters), and solve your inequality for w .

OpenStudy (anonymous):

perimeter:p length:l width:w p=2l+2w l=31 2l=62 and p <=100 so 2l+2w=62+2w<=100 so you gotta solve w for 62+2w<=100